OpenStudy (anonymous):

. Which of the following is the equation of a line that passes through (-2, -1) and (-4, -3)?

OpenStudy (swissgirl):

The slope is change of y over change of x (-1-(-3))/(-2-(-4) = (-1+3)/(-2+4) = 2/2=1 so the slope of the line is 1 We also know that an equation of a line is y=mx+b where m is the slope and b is the y intercept

OpenStudy (swissgirl):

So we have 2 diff points here so lets plug in (-4,-3) into the equation -3=(1)(-4)+b -3=-4+b 1=b so the line of equation that passes thru the points is y=x+1